# Design and Development of Portable RPC-Based Cosmic Muon Tracker

### v2

@@ Abstract
Primary cosmic rays when interact with our atmosphere, produce a cascade of lighter secondary particles namely pion, kaon, neutrons, muons, electrons, positrons and neutrinos. Muons are one of the most abundant and easily detectable particles at the ground surface using a large variety of particle detectors. Resistive Plate Chambers (RPCs) of in dimension were developed to be used in large scale as the active detector elements in the Iron Calorimeter (ICAL) which was planned to be built by the India-based Neutrino Observatory (INO). As a spin-off of this work, a portable stack of eight, one square foot RPC detectors is developed named as Cosmic Muon Tracker (CMT). It could be used to conduct small-scale particle detector experiments as well as training Students. We will discuss design, integration, characterisation and some of the applications of this detector in this paper.
@@ 1 Introduction
When high energy cosmic rays - primarily the protons, hit earth’s atmosphere, they produce showers of secondary particles comprising mostly pions. The pions further decay and produce muons and neutrinos. This shower of particles initially increases and then reduces in number as they travel towards the earth’s surface. As a result, the average flux of muons is about 1 per square centimeter per minute at mean sea level. The muon flux is a function of altitude [1] and their angular momentum depends on the earth’s magnetic field at that location. The wide spectrum of energy and angular distributions makes muons a useful radiation probe for various scientific as well as societal applications such as medical imaging, tomography, structural damage of a bridge/building and various particle physics experiments. The charged cosmic ray muons can be detected using a variety of detectors such as scintillators, Resistive Plate Chambers (RPCs), proportional counters, etc. In general, multiple large area planar detectors arranged in a stacked geometry are used to track muons. RPCs are gaseous detectors [2] which are widely used in particle physics experiments due to their excellent timing and large area coverage, besides ease of fabrication and low cost.
@@
The Iron CALorimeter (ICAL) experiment, proposed by the India-based Neutrino Observatory (INO) collaboration [3], aims to track and study atmospheric neutrinos using a 50-kiloton iron calorimeter. The detector comprises 28,800 Resistive Plate Chambers (RPCs), each measuring , sandwiched between 56 mm thick iron plates to track charged particles produced by neutrino interactions. Each RPC is equipped with a Front-End Data Acquisition module, known as RPC-DAQ, which records the position and timing information of particle trajectories. The collected data is transmitted to back-end servers via an Ethernet interface, with each RPC-DAQ remotely controllable through a dedicated IP address. Event registration and synchronization are managed using a Real-Time Clock system. Based on the detector and instrumentation R&D for the ICAL experiment, a portable detector stack of about one cubic foot in size, capable of tracking cosmic muons in real time is developed. The stack is made up of eight RPCs of about 1 ft × 1 ft in dimension which are filled with a special gas mixture and sealed. Eight Equally segmented signal strips on the top and bottom of the RPC gap of all eight RPCs are readout and processed. Strips on top and bottom layer placed orthogonally and marked as X- and Y-strips. Typically, cosmic ray muons passing through the stack leave tracks in all the eight RPCs and induce signals on the pickup strips of all the planes, which are in the path of the muon trajectory. LEDs - one per strip, mounted on the front-end boards of X- and Y-planes of each RPC in the stack, display the tracks of the cosmic ray muons in real-time as shown in Figure 1. The strip hit information is also stored in an FPGA-based digital Data Acquisition (DAQ) module, which is connected to a computer through Ethernet. The DAQ module also produces the event trigger whenever the input signals match the user-defined trigger criteria. This trigger signal is used to acquire the data from all the strips/channels. The data acquired is transferred to the server and is also displayed as an event track using the on-detector LED matrix. The live display of cosmic muon events can also be ported onto the internet for viewing from anywhere. The low voltages required for powering electronics as well as to generate the high voltage required to bias the RPCs are all locally generated from mains power, thus making this tracker highly portable. The paper describes design and fabrication of the detectors, electronics, trigger and data acquisition systems, performance of the stack as well as its other potential applications.
2 Detector Geometry
As shown in Figure 2, the Cosmic Muon Tracker (CMT) consists of 8 RPCs stacked vertically and each separated by . Each layer is supported by thick aluminum honeycomb panels with C-type aluminum channels providing vertical support. Each RPC is centered in its layer using small L-shaped aluminum brackets. The overall size of the stack is with each RPC detector measuring Out of area of the RPC, only is the active detector surface for tracking muons. The X- and Y-sides of the stack are dedicated to the front-end modules and LED displays while the back of the X-side is used for cabling the high voltage distribution. A manifold made of PVC pipes used for gas distribution is mounted on the rear of the Y-side. The top surface of the stack is used for accommodating the data acquisition (DAQ) module along with low and high voltage modules. High-density differential signal cables connect each front-end module to the DAQ module and the low-voltage supply for each front-end is routed through the empty slots of the C-type channels from top to bottom of the stack. Transparent acrylic sheets are used to cover all sides of the stack to protect the front-end modules and RPCs.
3 Working principle of the RPC detector
RPCs consist of two thick parallel glass plates with a bulk resistivity of about . A semi-conductive paint is coated on one side of the glass plates. These plates are mounted parallel to each other - with uncoated sides facing each other, with a gap of between them which was maintained using poly-carbonate button-shaped spacers. The volume thus formed is sealed using 3M DP-190 glue on all four sides. This assembly is called the RPC gas gap. A bias voltage of about is applied across the gas gap using HV contacts pasted onto the conductive coating surfaces. Signal pickup is achieved using copper strip-based honeycomb panels placed on either side of the gas gap in an orthogonal configuration. Thin Mylar sheets are used to isolate the pickup strips from the conductive coating carrying high voltage.
The RPC gas gap is fitted with gas inlet and outlets to flow the ionization gas mixture consisting of R-134a (95%), isobutane(4.2%) and SF6(0.3%). through the chamber. Main components of an RPC detector are shown in Figure 3. RPC operates on the principle of ionization as shown in the Figure 4. A uniform electric field is created between the two plates inside the gas volume by applying a bias voltage across the chamber. When a charged particle (such as a cosmic ray muon) passes through the chamber, it ionizes the gas and thus generates free electrons and positive ions. The electrons are accelerated towards the anode due to the electric field, triggering a series of secondary ionizations on the way and thus leading to an avalanche.
@@ 3 Working principle of the RPC detector
This avalanche induces an electric signal on the copper strips attached to the external pickup panels. Each side of the RPC is divided into eight copper strips, each of width and a pitch of . The resistive nature of the glass plates ensures that the discharge is localized, and prevents it from spreading across the entire plate. The pickup strips are appropriately terminated to readout the signal from one end only. The signals are then transmitted through coaxial cables to the front-end electronics boards for further processing.
@@ 4 RPC Production and Testing
Before assembling the cosmic muon tracker, each RPC was fabricated and quality checked [4] with assistance from summer project students who were invited from various institutions across India. The students were assigned tasks such as measuring the surface resistivity of the conductive coating, assembling the RPC, leak rate estimation of the gas gaps, performing V-I characteristics and final characterization of the detectors.
@@ 4.1 Surface resistivity measurement
In order to ensure uniform distribution of the bias voltage and thus electric field across the chamber, the conductive coating must have uniform surface resistivity [5]. This is measured using a specially developed contact jig and a multi-meter. The jig consists of two copper coated brass pieces separated by a plastic spacer of the same length as that of the brass pieces. The enclosed paint surface by the jig thus forms a square. As shown in Figure 5 the jig is placed in 2 different orientations – called “H” and “I” in a matrix on the painted surface.
Figure 6 shows the surface resistivity measurements for an RPC in both orientations on top and bottom surfaces. Similar measurements were recorded for all the eight RPCs. Most of the painted glasses showed a uniform surface resistivity of about . It may be noted that the quantity measured by the multi-meter is resistance, while it actually represents a quantity called “resistance per square” due to construction of the jig as described above.
@@ 4.2 Leak test of RPC Gaps
The fabricated gas gaps must be gas-tight else the performance of the RPC could be compromised due to the contaminated gas mixture. To ensure that the RPCs are leak free, leak test measurements [6] were conducted using an Arduino-based system. The outlet of the RPC was connected to a pressure sensor linked to a NodeMCU microcontroller Wi-Fi module which transmits the pressure readings inside the chamber directly to a server. The inlet of the RPC was connected to a gas supply with a needle valve to finely adjust the pressure filling the chamber with nitrogen gas.
Another NodeMCU based pressure sensor module was used to measure the atmospheric pressure readings and transmitted to the server. A Python script was used to store the data and display the pressure difference over time. The pressure inside the RPC was increased above the atmospheric pressure and tightly sealed. The pressure inside the RPC and pressure difference are continuously monitored and logged at least for 5 hours. If the pressure remained stable within its initial range, then the chamber was deemed to be leak-free and is ready for further fabrication. The temperature inside the chamber was also plotted as a function of time.
@@ 4.3 I-V Characterization
Figure 8 shows the equivalent circuit of an RPC gas gap. At lower voltages, the current flows from the top plate to the bottom plate through the spacers with the resistance denoted as . When the RPC attains the operating potential, an electric field is established between the top and bottom plates and the current begins to flow through the gas medium whose resistance is represented as .
For the I-V characteristics test, the RPCs were connected to a differential high voltage supply and the voltage was gradually increased from to in steps of for each polarity. Channel A corresponds to positive polarity and channel B to negative polarity, meaning that a total potential difference of is applied when both channels reach a voltage of each. Current readings were recorded at each step after allowing the currents to stabilize. Figure 9 shows the I-V characteristics plots of both polarities of all the 8 RPCs used in the CMT.
@@ 4.4 RPC Detector Characterization
Figure 12 shows the noise rate of one of the RPC channels. It is essential to ensure that the noise rate of the RPC is minimal before conducting this measurement. Noise rate is one of the basic health monitoring parameters of the RPC. Noise rate of an RPC must be stable for proper tracking of muons through the RPC stack.
5 Stacking of CMT
After characterizing all eight RPC detectors, they were prepared for assembly into the CMT stack. Each RPC was placed on an aluminium honeycomb tray for rigidity and reduced weight. The detectors were fitted to the base plate using custom-fabricated “Z” clamps. Each of the eight layers was then mounted sequentially onto a rigid aluminium frame made from extruded sections as shown in Figure 13. Power cables for all the front-end (FE) boards were routed through the slot in the extruded section. Aluminium mounts for the FE boards on the X- and Y-sides were also fabricated to ensure rigid mounting. The complete stack of RPC layers was equipped with all necessary electronic components including FE boards, power supply boards, RPC data acquisition systems and gas supply manifolds. Cable harness and routing were done with all components connected to a common ground point using wires and copper strips to ensure a reliable ground reference. Signal cables were properly shielded to support high speed signal transmission and reduced noise.
Gas inlet of each RPC is connected in cascade to support flow of gas in all the RPCs uniformly. The pressure inside the RPC chambers were stable over time due to tight sealing of the gas inlets and outlets. The high voltage required for operation of the RPC is generated using an onboard power supply, thus enabling the CMT to be operated directly from an A.C. mains supply at any location.
@@ 6 Trigger and Data Acquisition System
To ensure the portability of the tracker, all the electronics comprising of Data Acquisition Module (DAQ), 16 Front-end circuits (FE), Low Voltage-High voltage (LVHV) module, power lines, signal cables and gas manifolds are harnessed carefully within the tracker geometry. The DAQ, LVHV and power supply modules are placed on top of the stack as shown in Figure 14.
The FE Boards on X- and Y-side are well interfaced with the FPGA-based Data Acquisition module named RPC-DAQ [7] shown in the Figure 15, which was originally developed as a digital Front-End for RPCs for the ICAL experiment. RPC-DAQ is capable of handling 128 inputs covering 64 channels from X-side and 64 channels from Y-side.
When a charged particle traverses the CMT, the RPCs strips in each layer and which are in the path of muon trajectory produce induced signals in the X- and Y-sides. The FE modules shown in Figure 16, in each layer amplify these signals and discriminate them above a fixed threshold. The resultant LVDS (Low Voltage Differential Signaling) logic signals are then sent to the RPC-DAQ module placed on top of the CMT unit.
These signals are processed within a few nanoseconds and the muon track is identified by a coincidence hardware logic shown in Figure 17. On successful coincidence – indicating passage of a cosmic muon through the stack, a trigger signal is generated to latch the event track hit information responsible for the trigger. This latched event is then sent back to the FEs and LEDs of corresponding RPC strips are lit, thus indicating live the cosmic track. This way, every muon generated track is displayed by the corresponding LEDs. Until the next trigger forms, the previous latched state is maintained in the LEDs of all layers.
Also, all the incoming signals to the FPGA are sent to a High-Performance Time to Digital Converter (HPTDC) [8] for time-stamping the input signals with resolution. This helps in tracking muons with accurate directional measurement. Also, a Real-Time Clock (RTC) is used to timestamp every muon event. Apart from LED tracking, the DAQ FPGA acquires position, timing and RTC information for every event and transfers the same to the computer using standard Ethernet protocols. Necessary hardware logics are implemented to continuously monitor the health of the detector. Both Event and health monitoring data are acquired by a "C" language-based Back-End (BE) server software. The BE server displays some basic information to the user from time to time and stores the acquired data. The BE server is equipped with features to configure various DAQ settings like trigger layers, RTC and channel masking at the start of the run. These features enhance the functionality of the user-friendly data acquisition for the tracker.

### v3

@@ Abstract
Primary cosmic rays when interact with the atmosphere, produce a cascade of lighter secondary particles namely pion, kaon, neutrons, muons, electrons, positrons and neutrinos. Muons are one of the most abundant and easily detectable particles at the ground surface using a large variety of particle detectors. Resistive Plate Chambers (RPCs) of in dimension were developed to be used in large scale as the active detector elements in the Iron Calorimeter (ICAL) which was planned to be built by the India-based Neutrino Observatory (INO). As a spin-off of this work, a portable stack of eight 26 cm 26 cm RPC detectors was developed, named the Cosmic Muon Tracker (CMT). It could be used to conduct small-scale particle detector experiments as well as training Students. We will discuss design, integration, characterization and some of the applications of this detector in this paper.
@@ 1 Introduction
High-energy cosmic rays, primarily protons, interact with Earth’s atmosphere, produce showers of secondary particles mostly pions. These pions further decay into muons and neutrinos. This shower of particles initially increases and then reduces in number as they travel towards the earth’s surface, resulting in a measurable muon flux of approximately one muon per square centimeter per minute at sea level. Muon flux varies with altitude [1] and their angular distribution is influenced by the Earth’s magnetic field at a given location. Due to their wide spectrum of energy, angular distributions and abundance, muons are used as radiation probe in diverse fields, including medical imaging, geophysical tomography, and high-energy physics. Several detector technologies have been used for muon detection, such as plastic scintillators [2], liquid scintillators [3], time projection chambers (TPCs) [4], semiconductor detectors detectors [5] and Resistive Plate Chambers (RPCs) [6]. Among them, RPC detectors offer advantages such as excellent time resolution, large-area coverage, ease of fabrication and cost-effectiveness. In general, multiple large area RPCs arranged in a stacked geometry are used to track muons.
@@
The Iron CALorimeter (ICAL) experiment, proposed by the India-based Neutrino Observatory (INO) collaboration [7], aims to track and study atmospheric neutrinos using a 50-kiloton iron calorimeter. The calorimeter comprises of 28,800 RPCs, each measuring , sandwiched between 56 mm thick iron plates to track charged particles produced by neutrino interactions. Each RPC is equipped with a Front-End Data Acquisition module, known as RPC-DAQ, which records the position and timing information of particle trajectories. The collected data is transmitted to back-end servers via an Ethernet interface, with each RPC-DAQ remotely controllable through a dedicated IP address. Event registration and synchronization are managed using a Real-Time Clock system. Based on the detector and instrumentation R&D for the ICAL experiment, we developed a compact, eight-layer RPC-based detector stack—named the Cosmic Muon Tracker (CMT)—that demonstrates real-time muon tracking and serves as a valuable tool for education and outreach. The detector stack consists of eight RPCs, each measuring approximately . These detectors are filled with a gas mixture and tightly packed for operational stability. Each RPC contains eight equally segmented signal strips on both the top and bottom surfaces. These strips are placed orthogonally and marked as X- and Y-strips, respectively, to enable two-dimensional position readout. Cosmic ray muons typically traverse multiple RPCs in the stack, inducing signals on the pickup strips intersected by their trajectories. LEDs—one per strip—mounted on the front-end boards of the X- and Y-planes illuminate in real time to display the muon tracks, as shown in Figure 1. To enable real-time visualization and recording of muon events, the CMT uses a compact FPGA-based Data Acquisition (DAQ) system. This system captures signals from all RPC layers and uses configurable logic to identify valid muon events. Upon detection, the corresponding strip hits are both logged and displayed using on-board LED matrices. The compact design, along with locally generated low and high voltage supplies from standard AC mains, makes the entire setup self-contained and highly portable. A more detailed description of the trigger logic and DAQ architecture is provided in Section 5. The paper describes design and fabrication of the detectors, electronics, trigger and data acquisition systems, performance of the stack as well as its other potential applications.
2 RPC Design Overview and Detection Principle
Each RPC comprises of two parallel 2 mm-thick glass plates with a bulk resistivity of approximately as measured at room temperature [8]. A semi-conductive paint is coated on one side of the glass plates. These plates are mounted parallel to each other - with uncoated sides facing each other. 2 mm thick polycarbonate button-shaped spacers are placed uniformly within the gap to maintain the separation. Edge and nozzle spacers are used to maintain uniform separation along the chamber edges, ensuring both mechanical stability and proper gas flow throughout the RPC volume. List of major components used to fabricate RPC is shown in Figure 2. 3M DP-190 epoxy adhesive was used to seal the chamber perimeter after assembling the plates and spacers. DP-190 was chosen for the INO ICAL RPCs for its non-reactive properties with most of the detector gases. Epoxy adhesive may exhibit mild out gassing, but no degradation in RPC performance (e.g., leakage current or efficiency) was observed during prior studies [6]. A bias voltage of about is applied across the outer coated surfaces, establishing a strong and uniform electric field inside the gap. The RPCs operate in avalanche mode with a gas mixture of R-134a (95.0%), iso-butane (4.2%), and SF6 (0.3%). R-134a serves as the primary ionizing medium, providing a high number of electron-ion pairs, iso-butane acts as a photon quencher to suppress secondary avalanches and stabilize operation, and SF6 functions as an electron quencher to minimize streamer formation. This composition is widely used in several large-scale RPC-based experiments [9] and has been time-tested in the long-term operation of INO-ICAL RPCs.
When a charged particle, such as a cosmic ray muon, passes through the gas-filled RPC, it can ionize gas molecules along its path, producing free electrons as well as positive and negative ions. The electrons are accelerated towards the anode by due to the electric field, generating a series of secondary ionizations and leading to an electron avalanche shown in Figure 3.
@@ 2 RPC Design Overview and Detection Principle
This avalanche induces an electric signal on the copper strips. Each side of the RPC is divided into eight copper strips, each of width and a pitch of . The resistive nature of the glass plates ensures that the discharge is localized, and prevents it from spreading across the entire plate. Thin Mylar sheets are used to isolate the pickup strips from the semi-conductive coating carrying high voltage. The pickup strips are appropriately terminated to readout the signal from one end only. The induced signals are transmitted to the front-end electronics for amplification, discrimination, and readout.
@@ 3 RPC Production and Testing
The RPCs used in the CMT were fabricated and tested in-house through a well-defined quality control process [10]. Undergraduate and graduate students from various institutions actively participated in the fabrication, quality checks, and characterization of RPCs as part of summer training and laboratory programs, gaining hands-on experience in detector construction. Students assisted with various aspects of the assembly and testing workflow. These included surface resistivity measurements of the resistive coating, leak testing of the gas gap, current-voltage (I–V) characterization, and efficiency measurements using cosmic ray muons. The testing process followed standardized protocols developed for large-scale RPC deployment in the INO-ICAL experiment. Details of the individual test procedures and results are presented in the following subsections.
@@ 3.1 Surface Resistivity Measurement
Uniform surface resistivity of the resistive paint coating on the outer glass surfaces is essential to ensure uniform voltage distribution and electric field uniformity across the RPC [11]. To measure this, a custom-built jig was used in conjunction with a digital multimeter. The jig consists of two parallel brass electrodes, separated by an insulating spacer, forming a well-defined square contact area on the painted surface. Measurements were performed on a grid covering each coated glass plate. The jig was placed in two orthogonal orientations as shown in Figure 4, labeled “H” and “I”, to account for any directional variation in coating uniformity.
Surface resistivity measurements across the glass surface were generally centered near , with observed variations up to a factor of 3 in some cases, as shown in Figure 5. These values remain within acceptable limits for stable RPC performance. It may be noted that the quantity measured by the multi-meter is resistance, while it actually represents a quantity called “resistance per square” due to construction of the jig as described above. These values are consistent with operational requirements for RPCs and comparable to those used in the INO-ICAL prototypes.
@@ 3.2 Leak test of RPC Gaps
The fabricated gas gaps must be gas-tight else the performance of the RPC could be compromised due to the contaminated gas mixture. To ensure that the RPCs are leak free, leak test measurements [12] were conducted using an Arduino-based system. The outlet of the RPC was connected to a pressure sensor linked to a NodeMCU microcontroller Wi-Fi module which transmits the pressure readings inside the chamber directly to a server. The inlet of the RPC was connected to a gas supply with a needle valve to finely adjust the pressure filling the chamber with nitrogen gas. Another NodeMCU based pressure sensor module was used to measure the atmospheric pressure readings and transmitted to the server. A Python script was used to store the data and display the pressure difference over time. The pressure inside the RPC was increased above the atmospheric pressure and tightly sealed. The pressure inside the RPC and pressure difference are continuously monitored and logged at least for 5 hours. If the pressure remained stable within its initial range, then the chamber was deemed to be leak-free and ready for further fabrication. The temperature inside the chamber was also plotted as a function of time.
@@ 3.3 I-V Characterization
Figure 7 shows the equivalent circuit of an RPC gas gap, where the leakage current at low bias is primarily attributed to conduction through the polycarbonate spacers, denoted as resistance between the plates. Epoxy used for sealing the chambers contributes negligibly to the overall current. No signs of surface leakage across epoxy-glass boundaries were observed during testing, as confirmed by stable current readings and visual inspection of glue lines. When the RPC attains the operating potential, an electric field is established between the top and bottom plates and the current begins to flow through the gas medium whose resistance is represented as .
For the I-V characteristics test, the total high voltage applied to the gap was increased from zero to 10000 kV in steps of 1000 V. Channel A corresponds to positive polarity and channel B to negative polarity, meaning that a total potential difference of is applied when both channels reach a voltage of each. Current readings were recorded at each step after allowing the currents to stabilize.
Figure 8 shows the current–voltage (IV) characteristics for all eight RPCs. While most chambers exhibit similar current behavior, RPC 1 and RPC 2 show slightly elevated currents across the bias range. The deviation in current observed at low voltages (2000–3000 V), where gas gain is negligible, is likely due to leakage currents across the spacers and the chamber borders. While surface irregularities or minor contamination during assembly may amplify these effects. However, these values remained well within safe operating limits and showed no sign of electrical breakdown or instability during extended testing.
@@ 3.4 RPC Detector Characterization
Noise rate is one of the basic health monitoring parameters of the RPC. Noise rate of an RPC must be stable for proper tracking of muons through the RPC stack. Figure 11 shows the noise rate of one of the RPC channels. The observed strip noise rates, typically in the range of 5–15 Hz, are consistent with expected values for similarly sized glass RPCs operated in avalanche mode. Similar noise distributions were observed across all eight chambers, with only a few strips requiring masking due to persistent high rates.
4 RPC integration into CMT Stack
The CMT consists of eight RPCs arranged vertically in a stack, with a uniform spacing of 35 mm between adjacent layers. Each RPC measures = , with the central = region defined as the active detection area. Figure 12 shows multiple views, including top, side, and a 3D rendering, of the complete eight-layer CMT assembly.
After individual testing and characterization, Each RPC is mounted on a 3 mm-thick aluminum honeycomb base-plate using custom-fabricated Z-shaped aluminum clamps, providing both structural rigidity and weight reduction. The detectors along with base-plates are placed within a custom-designed aluminum frame shown in Figure 13 using L-brackets and C-type vertical channels that provide mechanical alignment and support. The complete CMT stack houses all essential subsystems, including the RPCs, front-end electronics, power supply modules, data acquisition system, and gas manifolds. The overall size of the stack is .
Dedicated aluminum brackets were fabricated to mount the front-end boards on both X- and Y-sides of each layer. Power distribution cables for the front-end electronics were routed through designated slots in the frame. All components were connected to a common ground point using copper strips and grounding wires. Front-end electronics are mounted on the X- and Y-sides of the stack, while high-voltage cabling and gas manifolds are arranged along the rear and top. Power and signal cables are neatly routed through the vertical frame channels, and transparent acrylic panels enclose the sides of the stack to protect the RPCs and electronics while allowing visual access. Gas lines were connected in a cascading configuration to ensure uniform gas flow through each chamber. The internal pressure remained stable over time due to the high-quality sealing of inlet and outlet fittings. High voltage for RPC operation is generated locally using a compact onboard power module, allowing the entire tracker to operate directly from a standard power source without requiring external infrastructure.
@@ 5 Trigger and Data Acquisition System
The CMT employs a custom-designed, FPGA-based DAQ module named RPC-DAQ [13] originally developed for the INO-ICAL experiment shown in Figure 14. RPC-DAQ is capable of handling 128 inputs covering 64 channels from X-side and 64 channels from Y-side, interfaces with 16 front-end (FE) boards shown in Figure 15: 8 for the X-plane strips and 8 for the Y-plane strips.
When a charged particle traverses the CMT, it induces coincident signals in aligned strips of multiple RPC layers. Each FE board amplifies 8 analog inputs and outputs discriminated digital signals via Low Voltage Differential Signaling (LVDS) lines to the DAQ module. The DAQ logic shown in Figure 16 implements configurable trigger conditions based on coincidence between user-defined layers. When a valid trigger is detected - indicating passage of a cosmic muon through the stack, the hit pattern is latched and stored.
Simultaneously, the corresponding strip-level LEDs are activated, producing a real-time visual display of the muon track across the stack. This latched state is maintained until the next trigger event. The digitized hit data is timestamped using a High-Performance Time-to-Digital Converter (HPTDC) [14] with resolution, This helps in tracking muons with accurate directional measurement. Also, a Real-Time Clock (RTC) is used to timestamp every muon event. An Event data comprises of position, timing and RTC information of muon signals acquired from each layer. Necessary hardware logics are implemented to continuously monitor the health of the detector.
Event and monitoring data are transferred to a host PC via Ethernet using a custom server interface implemented in C. The server supports detector configuration (e.g., trigger layer selection, channel masking) and periodically displays health metrics such as strip noise rates and trigger counts. Live event data can also be streamed over the internet for remote monitoring or educational outreach. An integrated Low Voltage and High Voltage (LVHV) distribution module is used to power RPC-DAQ, FE boards and RPCs. The DAQ, LVHV and power supply modules are placed on top of the stack as shown in Figure 17.